OpenStudy (anonymous):

please help me factor the question: Prove: sin θ - sin θ•cos2 θ = sin3 θ. You must show all work.

OpenStudy (anonymous):

\[Prove: \sin \theta - \sin \theta.\cos ^{2}\theta = \sin ^{3} \theta\] \[= \sin \theta -(\sin \theta - \sin ^{3}\theta)\] \[\sin \theta - \sin \theta + \sin ^ {3} \theta = \sin^{3} \theta\]

OpenStudy (anonymous):

@silver321 :Your question's answered

OpenStudy (anonymous):

thank you :)